Mulitiply.
( r + 1/5)(r+1/3)

Answer by ewatrrr(24785)


Hi,
Mulitiplying
Note: 1/5 = 3/15 and 1/3 = 5/15
( r + 1/5)(r+1/3) = (r + 3/15)(r + 5/15)
r^2 + (8/15)r + 1/15
Note:SUM of the inner product(3/15)r and the outer product(5/15)r = (8/15)r